Solve
y + 2y = 8
y-2

Answer by checkley75(3666)

IF THIS IS THE PROBLEM (Y+2Y)/(Y-2)=8 THEN.
CROSS MULTIPLING WE GET:
8(Y-2)=Y+2Y
8Y-16=3Y
8Y-3Y=16
5Y=-16
Y=-16/5
Y=-3.2 ANSWER.